OpenStudy (anonymous):

Multiply. (x+2)(x-5) Simplify your answer.

OpenStudy (anonymous):

@jdoe0001 can you help me please

OpenStudy (sleepyjess):

Do you know the FOIL method?

OpenStudy (anonymous):

yes but its really confusing

OpenStudy (sleepyjess):

Ok, let's just take it one step at a time. We need to multiply to first term of each binomial together first.

OpenStudy (jdoe0001):

\(\large { (x+2)(x-5)\implies \begin{array}{llll} x+2 \\ \quad \\ \times x \\\hline\\ \end{array}+\begin{array}{llll} x+2 \\ \quad \\ \times -5 \\\hline\\ \end{array} }\) or use the FOIL method http://www.mathsisfun.com/definitions/foil-method.html

OpenStudy (sleepyjess):

*multiply the

OpenStudy (anonymous):

so x(x) = x^2

OpenStudy (anonymous):

I don't understand :-/

OpenStudy (sleepyjess):

Yep, first is x*x which is x^2

OpenStudy (sleepyjess):

Now the O represents the outside terms. So the terms on the front and back end.

OpenStudy (anonymous):

is it x^2 + 5x - 2x + 10?

OpenStudy (sleepyjess):

Close, you got + 5x correct, but for the last 2 terms, it should be 2*x and 2*-5

OpenStudy (anonymous):

x^2+3x-5?

OpenStudy (sleepyjess):

What is 2*x

OpenStudy (anonymous):

Yeah you're really close to the answer so first we multiply x*x = x^2 next you would multiply x*-5 = -5x then you would multiply 2*x= 2x and lastly you would multiply 2*-5= -10 SO then you would add up all the terms you found together: (x^2) + (-5x) + (2x) + (-10) Which equals (x^2) - 3x - 10 [Think of the distributive property] With a bit more practice it'll get easier and easier for you :)